1. What is the yield to maturity of a eighteight​-year, $10,000 bond with a 4.8​% coupon rate and semiannual coupons if this bond is currently trading for a price of $8,740​?

2.What is the present value​ (PV) of $100,000 received five years from​ now, assuming the interest rate is 99​% per​ year?

3. What is the coupon rate of a eighteight​-year, $10,000 bond with semiannual coupons and a price of $8,637.58​, if it has a yield to maturity of 7.4​%?

4. How much will the coupon payments be of a 25​-year $10,000 bond with a 10​% coupon rate and semiannual​ payments?
Answer #1

1. Price =8740
Par Value =10000
Coupon =4.8%*10000/2 =240
Number of periods =8*2 =16
YTM using excel formula =2*RATE(16,240,-8740,10000) =6.87%

2. FV =100000
Number of years =5
Rate =99%
PV =FV/(1+r)^n =100000/(1+99%)^5 =3204.31